OpenStudy (anonymous):

The sum of a number and its square is 72. Find the number.

OpenStudy (anonymous):

Let the number be x therefore \[\huge x^2 +x = 72 \rightarrow x^2 +x -72=0\] \[\huge \rightarrow x^2 +9x-8x -72=0\] \[\huge \rightarrow x(x +9)-8(x +9)=0\] \[\huge \rightarrow (x +9)(x-8)=0\] \[\huge \rightarrow (x +9)=0 or (x-8)=0\] \[\huge x=-9 or x=8\] Hence the required numbers are -9 or 8

OpenStudy (anonymous):

@gg0824